missrosejensen:

Which of the following equations has a solution of -50? = -25 -2r = 50 b - (20) = -30 n + (-16) = 65

Shadow:

The first choice looks like the equation got chopped in half

missrosejensen:

yeah, the first one is x over 2 =-24

Shadow:

Can you take a screenshot or just write it out.

Shadow:

Can't log into your student account, that's a bit too far.

missrosejensen:

Shadow:

\[\frac{ x }{2 } = -25\] The only way to isolate x here is to multiply both sides by 2 \[2 \times \frac{ x }{2 } = -25 \times 2\] \[x = -50\]

missrosejensen:

thanks again